OpenStudy (anonymous):

Simplify. x(4-3x)+4(x^2+1)/x^2-4

OpenStudy (anonymous):

\[\frac{ x(4-3x)+4(x ^{2}+1) }{ x ^{2}-4 }\]

OpenStudy (anonymous):

@satellite73 I'm so sorry :/ I've been having yu hold my hand thru all these problems

OpenStudy (anonymous):

donkey work if ever i saw it http://www.wolframalpha.com/input/?i=%28x%284-3x%29%2B4%28x^2%2B1%29%29%2F%28x^2-4%29

OpenStudy (anonymous):

but we can still work it out step by step first expand then combine like terms then factor then cancel a bunch ready ?

OpenStudy (anonymous):

Ready

OpenStudy (anonymous):

\[\frac{ x(4-3x)+4(x ^{2}+1) }{ x ^{2}-4 }\] \[=\frac{ 4x-12x^2+4x^2+4 }{ x ^{2}-4 }\] is a start

OpenStudy (anonymous):

The numerator, wouldn't it just be \[4x-3x ^{2}\] instead of \[4x-12x ^{2}\]

OpenStudy (anonymous):

yeah i screwd it up

OpenStudy (anonymous):

\[=\frac{ 4x-3x^2+4x^2+4 }{ x ^{2}-4 }\]

OpenStudy (anonymous):

so you get \[\frac{x^2+4x+4}{x^2-4}\] factor the top on bottom you good with that?

OpenStudy (anonymous):

both factor very easily, even i can do it

OpenStudy (anonymous):

Haha :) give me a minute

OpenStudy (anonymous):

no do it instantly in your head!

OpenStudy (anonymous):

\[\frac{ (x+2)(x+2) }{ (x+2)(x-2) }\]

OpenStudy (anonymous):

..I'm a slow math person shush

OpenStudy (anonymous):

\[\frac{ (x+2) }{ (x-2) }\]

OpenStudy (anonymous):

if you can do that other factoring you can certainlyi recognize a perfect square \(x^2+4x+4=(x+2)^2\) and the difference of two square \(x^2-4=(x+2)(x-2)\) when you see them

OpenStudy (anonymous):

and yes you get what you said too bad though

OpenStudy (anonymous):

Too bad?

OpenStudy (anonymous):

because problems like these lead you to believe that whenever you see a fraction like that (in some future real math class) that you will be able to factor and cancel in reality you can rarely do it, but since you have been trained to think that everything goes, you may be tempted to cancel even when you cannot

OpenStudy (anonymous):

just saying if watch out in the future guard against cancelling incorrectly it is a common mistake

OpenStudy (anonymous):

Alrighty, I'll keep that in mind. Thanks! :)

OpenStudy (anonymous):

you done?